# Adaptive-TMLE for the Average Treatment Effect based on Randomized Controlled Trial Augmented with Real-World Data

### v1

@@ 2.2 Statistical model
Assumption A1 is a causal assumption and is non-testable, assumptions A2 and A3 are statistical assumptions. Note that assumptions A1 and A2 are satisfied in an RCT (or a well-designed observational study). Importantly, we do not make the assumption that $E(Y_{a}\mid S=0,W)=E(Y_{a}\mid W)$, sometimes referred to as the “mean exchangeability over $S$[dahabreh_generalizing_2019]. In other words, for the combined study, $A$ might not be conditionally independent of $Y_{0},Y_{1}$, given $W$, due to the bias introduced from the external data. Specifically, we are concerned that $E_{W}E(Y_{1}-Y_{0}\mid W)\not=E_{W}E(Y\mid W,A=1)-E_{W}E(Y\mid W,A=0)$, due to $E_{W}E(Y_{1}-Y_{0}\mid W,S=0)\not=E_{W}E(Y\mid S=0,W,A=1)-E_{W}E(Y\mid S=0,W,A=0)$. Beyond these three assumptions, we will make additional assumptions on $g_{0}(1\mid S,W)\equiv P(A=1\mid S,W).$ In particular, if $S=1$ corresponds with an RCT, then $g_{0}(1\mid S=1,W)$ would just be the randomization probability; If the $S=1$-study is not an RCT, we might still be able to make a conditional independence assumption $g_{0}(A\mid 1,W)=g_{0}(A\mid 1,W_{1})$ for some subset $W_{1}$ of $W$. In some cases, we might also be able to assume that $S$ is independent of $W$, i.e. $P_{W\mid S=1}=P_{W\mid S=0}.$ This independence could be achieved by sampling the external subjects from the same target population as in the $S=1$-study. As we will see later in the identification step, for our target parameter it is crucial that the support of $P_{W\mid S=0}$ is included in the support of $P_{W\mid S=1}$ so that assumption A3 holds. In this article we will not make assumptions on the joint distribution of $(S,W)$ beyond assumption A3, but our results are not hard to generalize to the case that we assume $W\perp S$. In the important case that we augment an RCT with external controls only we have that $g_{0}(1\mid 0,W)=0$, so that the only purpose of the $S=0$-study is to augment the control arm of the RCT. To briefly summarize, we do not make any additional assumptions beyond the same set of assumptions for a standard RCT. The only additional assumption we require is A3 and could be made plausible in the selection of the external subjects.
@@
$p(s,w,a,y)=p_{S}(s)p_{W}(w\mid s)g_{A}(a\mid s,w)q_{Y}(y\mid s,w,a),$
where $q_{Y}(y\mid s,w,a)\equiv p_{Y}(y\mid s,w,a).$ Our statistical model only makes assumptions on $g_{A}$ and leaves the other factors in the likelihood unspecified.
@@ 2.3 Target causal parameters
We could consider two candidate target parameters. The first one is
@@
Note that this parameter measures the conditional treatment effect in the RCT, while it takes an average with respect to the combined covariate distribution $p_{W}=p_{W\mid S=1}p_{S=1}+p_{W\mid S=0}p_{S=0}$. Alternatively, we could take the average with respect to the RCT covariate distribution $p_{W\mid S=1}$, in which case we have the target parameter given by
$\Psi^{F}_{2}(P_{O,U})=E(Y_{1}-Y_{0}\mid S=1)=E_{W}[E(Y_{1}-Y_{0}\mid S=1,W)\mid S=1].$
In the following subsections, we will discuss the their identification results and compare their efficient influence functions.
@@ 2.4 Statistical estimand
Note that this estimand is only well-defined if $\min_{a\in\{0,1\}}P(S=1,W=w,A=a)>0$ for $P_{W}$-a.e. So, if the $S=0$-study has a covariate distribution with a support not included in the support of $P_{W\mid S=1}$, then the estimand is not well-defined. Therefore, we need assumption A2. As we will see in the next subsection, the efficient influence function of $\Psi$ indeed involves inverse weighting by $P(S=1\mid W)$. Similarly, $\Psi^{F}_{2}$ is identified by
@@
We could construct an efficient estimator of $\Psi:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$. However, it would still lack power since it would only be using the $S=1$-observations for learning the conditional treatment effect, and similarly for $\Psi_{2}$. This point is further discussed in the next subsection by analyzing and comparing the nonparametric efficiency bounds of $\Psi$ and $\Psi_{2}.$ Therefore, we instead pursue a finite sample super-efficient estimator through adaptive-TMLE whose gain in efficiency is adapted to the underlying unknown (but learnable) complexity of $P_{0}$.
@@ 2.5 Efficient influence functions of the target causal parameters
$D_{\Psi,P}(O)=\bar{Q}_{P}(1,W,1)-\bar{Q}_{P}(1,W,0)-\Psi(P)+\frac{S}{P(S=1\mid W)}\cdot\frac{2A-1}{g_{P}(A\mid 1,W)}(Y-\bar{Q}_{P}(S,W,A)).$
This is also the canonical gradient in the statistical model that assumes additionally that $S$ is independent of $W$.
The canonical gradient of $\Psi_{2}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$ at $P$ is given by:
$D_{\Psi_{2},P}(O)=\frac{S}{P(S=1)}(\bar{Q}_{P}(1,W,1)-\bar{Q}_{P}(1,W,0)-\Psi_{2}(P))+\frac{S}{P(S=1)}\cdot\frac{2A-1}{g_{P}(A\mid 1,W)}(Y-\bar{Q}_{P}(S,W,A)).$
@@
$PD^{2}_{\Psi_{2},P}(O)$ $=E_{P}\left[\frac{P(S=1\mid W)}{P^{2}(S=1)}(\bar{Q}_{P}(1,W,1)-\bar{Q}_{P}(1,W,0)-\Psi_{2}(P))^{2}\right]$
@@
Thus, the variance of the $W$-component of $D_{\Psi,P}$ is a factor of $\sim 1/P(S=1)$ smaller than the variance of the $W$-component of $D_{\Psi_{2},P}$. However, the variance of the $Y$-component of $D_{\Psi,P}$ involves the factor $1/P(S=1\mid W)$ versus the factor $\sim 1/P(S=1)$ in $D_{\Psi_{2},P,Y}$. Therefore, if $S$ is independent of $W$, then it follows that the variance of $D_{\Psi,P}$ is smaller than the variance of $D_{\Psi_{2},P}$, due to a significantly smaller variance of its $W$-component, while having identical $Y$-components. On the other hand, if $S$ is highly dependent on $W$, then the inverse weighting $1/P(S=1\mid W)$ could easily cause the variance of the $Y$-component of $D_{\Psi,P}$ to be significantly larger than the variance of the $Y$-component of $D_{\Psi_{2},P}$. Generally speaking, especially when $P(A=1\mid S=1,W)$ depends on $W$, the variance of the $Y$ component dominates the variance of the $W$ component. Therefore without controlling the dependence of $S$ and $W$, one could easily have that the variance of $D_{\Psi,P}$ is larger than the variance of $D_{\Psi_{2},P}$. Hence, if one wants to use $\Psi$ instead of $\Psi_{2}$, then one wants sample external subjects such that $S$ approximately independent of $W$. For this article, we will focus our attention on the target causal parameter $\Psi(P)$, which, as we argued above, would be the preferred choice if $P(S=1\mid W)\approx P(S=1)$. However, our results can be easily generalized to $\Psi_{2}(P)$.
@@
As mentioned earlier, without additional assumptions, one could construct an efficient estimator for $\Psi(P_{0})$. However, it may still lack power (we will also show this empirically through simulations in Section 8). On the other hand, if investigators are willing to make the assumption that $E(Y_{a}\mid S=0,W)=E(Y_{a}\mid W)$, then a more efficient target estimand would be the pooled-ATE estimand, given by
$\tilde{\Psi}(P_{0})\equiv E_{0}[E_{0}(Y\mid A=1,W)-E_{0}(Y\mid A=0,W)],$
where one simply pools the two studies. Under this assumption, we would be in an optimal scenario from a power perspective with $\Psi(P)=\tilde{\Psi}(P).$ However, this assumption may not hold in practice, for example, due to unmeasured confounding in the external data. This then motivates us to define the bias-estimand as
@@
In other words, $\Psi^{\#}$ is exactly the bias introduced by using $\tilde{\Psi}$ instead of $\Psi$ as the target. Then, it follows that we could write our target estimand as
@@
Now, we can view our target estimand as applying a bias correction $\Psi^{\#}$ to the (potentially biased) pooled-ATE estimand $\tilde{\Psi}$. For estimating $\tilde{\Psi}(P_{0})$, which is simply an average treatment effect on the pooled data, one could follow the recipe in [adml] to construct an A-TMLE for it. Ingredients for constructing it is detailed in Section 5. For now, let’s focus on the bias estimand. Since A-TMLE is a general framework for estimating conditional effects, we need to parameterize the bias estimand as a function involving conditional effects. The following lemma allows us to express $\Psi^{\#}$ in terms of the effect of $S$ on $Y$ conditioned on $W$ and $A$, and the distribution of $S$ conditioned on $W$ and $A$.
@@
Let
$\tau_{P}(W,A)$ $\equiv E_{P}(Y\mid S=1,W,A)-E_{P}(Y\mid S=0,W,A)$
@@
$=-P(S=0\mid W,A=0,)[E_{P}(Y\mid S=1,W,A=0)-E_{P}(Y\mid S=0,W,A=0)].$
Thus, $E_{W}E(Y\mid W,A=0)-E_{W}E(Y\mid S=1,W,A=0)=-E_{W}\Pi_{P}(0\mid W,0)\tau_{P}(W,0)$. Therefore, $\Psi^{\#}(P)=\tilde{\Psi}(P)-\Psi(P)=E_{W}[\Pi_{P}(0\mid W,0)\tau_{P}(W,0)-\Pi_{P}(0\mid W,1)\tau_{P}(W,1)]$. ∎
Lemma 2 shows that the bias estimand can be viewed as the expectation of a weighted combination of the conditional effect of the study indicator $S$ on $Y$ of the two treatment arms, where the weights are the probabilities of enrolling in the RCT of the two arms. Now, we have successfully write the bias estimand as a function of two conditional effects. In the next two sections, we will proceed to discuss the ingredients and steps to construct A-TMLEs for $\Psi^{\#}$ and $\tilde{\Psi}$, respectively.

### v2

@@ 2.2 Statistical model
Note that assumptions A1 and A2 are satisfied in an RCT (or a well-designed observational study without unmeasured confounding). Importantly, we do not make the assumption that $E(Y_{a}\mid S=0,W)=E(Y_{a}\mid W)$, sometimes referred to as the “mean exchangeability over $S$[dahabreh_generalizing_2019]. In other words, for the combined study, the treatment might not be conditionally independent of the potential outcomes given baseline covariates, due to the potential bias introduced by the external data, such as unmeasured confounding. Specifically, we are concerned that $E_{W}E(Y_{1}-Y_{0}\mid W)\not=E_{W}E(Y\mid W,A=1)-E_{W}E(Y\mid W,A=0)$, due to $E_{W}E(Y_{1}-Y_{0}\mid W,S=0)\not=E_{W}E(Y\mid S=0,W,A=1)-E_{W}E(Y\mid S=0,W,A=0)$. Beyond these three assumptions, we have additional knowledge on $P_{0}(A=1\mid S,W).$ In particular, if $S=1$ corresponds to an RCT, then $P_{0}(A=1\mid S=1,W)$ would just be the randomization probability; If the $S=1$-study is not an RCT, we might still be able to make a conditional independence assumption $P_{0}(A=1\mid S=1,W)=P_{0}(A=1\mid S=1,W=W_{1})$ for some subset $W_{1}$ of $W$. In some cases, we might also be able to assume that $S$ is independent of $W$, i.e. $P_{W\mid S=1}=P_{W\mid S=0}.$ This independence could be achieved by sampling the external subjects from the same target population as in the $S=1$-study. As we will see later in the identification step, for our target parameter it is crucial that the support of $P_{W\mid S=0}$ is included in the support of $P_{W\mid S=1}$ so that assumption A3 holds. In this article we will not make assumptions on the joint distribution of $(S,W)$ beyond assumption A3, but our results are not hard to generalize to the case that we assume $W\perp S$. In the important case that we augment an RCT with external controls only we have that $P_{0}(A=1\mid S=0,W)=0$, so that the only purpose of the $S=0$-study is to augment the control arm of the RCT. To briefly summarize, we only rely on assumptions known to hold for a well-designed RCT. The only additional assumption we require is A3 and could be made plausible in the selection of the external subjects.
@@
$p(s,w,a,y)=p_{S}(s)p_{W}(w\mid s)p_{A}(a\mid s,w)p_{Y}(y\mid s,w,a).$
Our statistical model only leverages knowledge on $p_{A}$ and leaves the forms of other factors in the likelihood unspecified.
@@ 2.3 Target causal parameters
We could consider two candidate target parameters. The first one is the covariate-pooled ATE:
@@
Note that this parameter measures the conditional treatment effect in the RCT, while it takes an average with respect to the RCT and external RWD combined covariate distribution $p_{W}=p_{W\mid S=1}p_{S=1}+p_{W\mid S=0}p_{S=0}$. Alternatively, we could take the average with respect to the RCT-only covariate distribution $p_{W\mid S=1}$, in which case we have the RCT-only ATE target parameter:
$\Psi^{F}_{2}(P_{O,U})=E_{W}[E(Y_{1}-Y_{0}\mid S=1,W)\mid S=1].$
In the following subsections, we will discuss their identification results and compare their efficient influence functions.
@@ 2.4 Statistical estimand
Note that this estimand is only well-defined if $\min_{a\in\{0,1\}}P(S=1,W=w,A=a)>0$ for $P_{W}$-a.e. So, if the $S=0$-study has a covariate distribution with a support not included in the support of $P_{W\mid S=1}$, then the estimand is not well-defined. This explains why assumption A3 is needed. As we will see in the next subsection, the efficient influence function of $\Psi$ indeed involves inverse weighting by $P(S=1\mid W)$. Similarly, $\Psi^{F}_{2}$ is identified by
@@
We could construct an efficient estimator of $\Psi:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$. However, it may still lack power since it would only be using the $S=1$-observations for learning the conditional treatment effect, and similarly for $\Psi_{2}$. This point is further discussed in the next subsection by analyzing and comparing the nonparametric efficiency bounds of $\Psi$ and $\Psi_{2}.$ Therefore, we instead pursue a finite sample super-efficient estimator through adaptive-TMLE whose gain in efficiency is adapted to the complexity of the unknown (but learnable) $P_{0}$.
@@ 2.5 Efficient influence functions of the target causal parameters
$D_{\Psi,P}(O)=Q_{P}(1,W,1)-Q_{P}(1,W,0)-\Psi(P)+\frac{S}{P(S=1\mid W)}\cdot\frac{2A-1}{g_{P}(A\mid 1,W)}(Y-Q_{P}(S,W,A)).$
This is also the canonical gradient in the statistical model that assumes additionally that $S$ is independent of $W$. The canonical gradient of $\Psi_{2}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$ at $P$ is given by:
$D_{\Psi_{2},P}(O)=\frac{S}{P(S=1)}(Q_{P}(1,W,1)-Q_{P}(1,W,0)-\Psi_{2}(P))+\frac{S}{P(S=1)}\cdot\frac{2A-1}{g_{P}(A\mid 1,W)}(Y-Q_{P}(S,W,A)).$
@@
$PD^{2}_{\Psi_{2},P}(O)$ $=E_{P}\left[\frac{P(S=1\mid W)}{P^{2}(S=1)}(Q_{P}(1,W,1)-Q_{P}(1,W,0)-\Psi_{2}(P))^{2}\right]$
@@
Thus, the variance of the $W$-component of $D_{\Psi,P}$ is a factor of $\sim 1/P(S=1)$ smaller than the variance of the $W$-component of $D_{\Psi_{2},P}$. However, the variance of the $Y$-component of $D_{\Psi,P}$ involves the factor $1/P(S=1\mid W)$ versus the factor $\sim 1/P(S=1)$ in $D_{\Psi_{2},P,Y}$. Therefore, if $S$ is independent of $W$, then it follows that the variance of $D_{\Psi,P}$ is smaller than the variance of $D_{\Psi_{2},P}$, due to a significantly smaller variance of its $W$-component, while having identical $Y$-components. On the other hand, if $S$ is highly dependent on $W$, then the inverse weighting $1/P(S=1\mid W)$ could easily cause the variance of the $Y$-component of $D_{\Psi,P}$ to be significantly larger than the variance of the $Y$-component of $D_{\Psi_{2},P}$. Generally speaking, especially when $P(S=1\mid W)$ depends on $W$, the variance of the $Y$ component dominates the variance of the $W$ component. Therefore, without controlling the dependence of $S$ and $W$, one could easily have that the variance of $D_{\Psi,P}$ is larger than the variance of $D_{\Psi_{2},P}$. If the goal of data integration is to maximize statistical efficiency and $\Psi$ is preferred over $\Psi_{2}$, it is important to sample external subjects such that $S$ is approximately independent of $W$. On the other hand, if the aim is to enhance the generalizability of trial findings, inducing dependence between $S$ and $W$ may be reasonable, as it aligns the covariate distribution in the pooled population more closely with that of the trial’s intended target population. For example, if certain patient subgroups are underrepresented in the trial, increasing the likelihood of sampling these patients in the external cohort may be appropriate to enhance the interpretation of estimates. For this article, given that our data application prioritizes efficiency gains, we will focus our attention on the target causal parameter $\Psi(P)$, which, as we argued above, would be the preferred choice if $P(S=1\mid W)\approx P(S=1)$. However, our results can be easily generalized to $\Psi_{2}(P)$.
@@
One could construct an A-TMLE directly for $\Psi(P_{0})$. This requires data-adaptively learning a working model for the conditional average treatment effect (CATE) within the RCT and then constructing a TMLE for the projection estimand implied by this working model [adml]. However, in our data augmentation problem context, such an A-TMLE would not utilize the external data for estimating the outcome regression, as the CATE function is conditioned on $S=1$, i.e., the RCT. This motivates us to consider the following alternative approach. First, we use A-TMLE to estimate a target estimand that prioritizes statistical power by fully utilizing the pooled data from the RCT and external RWD, while acknowledging that it may be biased. Next, we apply A-TMLE again to estimate and correct for this bias. In our setting, a potentially more efficient estimand to target is the pooled-ATE estimand, given by:
$\tilde{\Psi}(P_{0})\equiv E_{0}[E_{0}(Y\mid A=1,W)-E_{0}(Y\mid A=0,W)].$
However, as we argued in Section 2, making this estimand equal to $\Psi(P_{0})$ would rely on identification assumptions that may not hold in many scenarios. Therefore, instead of assuming $\tilde{\Psi}=\Psi$, we define the bias estimand:
@@
In other words, $\Psi^{\#}$ captures all the bias introduced by using $\tilde{\Psi}$ instead of $\Psi$ as the target. Then, it follows that we could write our target estimand as:
@@
Now, we can view our target estimand as applying a bias correction $\Psi^{\#}$ to the (potentially biased) pooled-ATE estimand $\tilde{\Psi}$. In order to apply the type of A-TMLE as presented in [adml], one needs to parameterize the target estimand as an average of conditional effects. Note that, $\tilde{\Psi}(P)$ is already an average of the conditional treatment effect, one could follow the approach outlined in [adml] to construct an A-TMLE for it. We now turn our attention to the bias estimand. Lemma 2 expresses $\Psi^{\#}$ in terms of the conditional average RCT-enrollment effect and the RCT enrollment mechanism.
@@
Let $\tau_{S,P}(W,A)$ and $\Pi_{P}(s\mid W,A)$ be the conditional average RCT-enrollment effect and the trial enrollment mechanism, where
$\tau_{S,P}(W,A)$ $\equiv E_{P}(Y\mid S=1,W,A)-E_{P}(Y\mid S=0,W,A)$
@@
$=-P(S=0\mid W,A=0)[E_{P}(Y\mid S=1,W,A=0)-E_{P}(Y\mid S=0,W,A=0)].$
Thus, $E_{W}E(Y\mid W,A=0)-E_{W}E(Y\mid S=1,W,A=0)=-E_{W}\Pi_{P}(0\mid W,0)\tau_{S,P}(W,0)$. Therefore, $\Psi^{\#}(P)=\tilde{\Psi}(P)-\Psi(P)=E_{W}[\Pi_{P}(0\mid W,0)\tau_{S,P}(W,0)-\Pi_{P}(0\mid W,1)\tau_{S,P}(W,1)]$. ∎
Now, we have parameterized the bias estimand as a weighted combination of two conditional effects, $\tau_{S,P}(W,0)$ and $\tau_{S,P}(W,1)$, where the weights are the probabilities of enrolling in the RCT of the two arms. This estimand resembles the ATE estimand and thereby enables a similar type of A-TMLE as for $\tilde{\Psi}$ with the role of CATE replaced by $\tau_{S,P}$.
Under the decomposition of the target estimand, the amount of efficiency gain from A-TMLE is driven by the complexities of $E(Y\mid W,A=1)-E(Y\mid W,A=0)$, the CATE function in the pooled data, and the bias function $E(Y\mid S=1,W,A)-E(Y\mid S=0,W,A)$. If reasonable efforts are made to ensure that the bias $\Psi^{\#}$ is small, a simple function might suffice to approximate the true (potentially fully nonparametric) bias function well, thereby offering efficiency gains. On the other hand, if the bias is large in magnitude but can be well-approximated by a parsimonious model, such as an intercept-only model, then we would also expect gains in efficiency. Intuitively, the decomposition strategy can be viewed as a general way of leveraging A-TMLE to achieve efficiency gains. The process begins by identifying an “easy-to-estimate” target, which, while potentially biased, is straightforward to estimate efficiently using all the available data. In our context, this would naturally be the pooled-ATE estimand, which can be estimated using the pooled data. Next, we parameterize the bias in terms of conditional effects and use A-TMLE to estimate the bias. The choice of the easy-to-estimate target should be guided by the specific problem setting and the nature of the resulting bias function.
In this section, we outline the key components and steps to construct an A-TMLE for the pooled-ATE estimand, $\tilde{\Psi}(P_{0})$. Since the pooled-ATE is the ATE on the pooled sample (RCT and external data combined), the estimation procedure mirrors that of the ATE example presented in [adml]. However, for completeness, we present it here within the context of our problem.
4.1 Semiparametric regression working model for the conditional average treatment effect
Given the fact that $\bar{Q}_{0}(W,A)\equiv E_{0}(Y\mid W,A)=\bar{Q}_{0}(W,0)+A\tau_{A,0}(W)$, a working model $\mathcal{T}_{A,w}$ for $\tau_{A,0}$ corresponds to a semiparametric regression working model $\bar{{\cal Q}}_{w}=\{\theta+A\tau_{A,\beta}:\beta,\theta\}$ for $\bar{Q}_{0}$, where $\theta$ is an unspecified function of $W$. This working model $\bar{{\cal Q}}_{w}$ further implies a corresponding working model ${\cal M}_{A,w}=\{P\in{\cal M}:\bar{Q}_{P}\in\bar{{\cal Q}}_{w}\}\subset{\cal M}$ for the observed data distribution $P_{0}$. Let $\tau_{A,\beta_{P}}\in{\cal T}_{A,w}$ be a projection of $\tau_{A,P}$ onto the working model ${\cal T}_{A,w}$ with respect to a specified loss function. This generally corresponds to mapping a $P\in{\cal M}$ into a projection $\Pi_{\mathcal{M}_{A,w}}(P)\in{\cal M}_{A,w}$. We recommend using the squared-error projection of $\bar{Q}_{P}$ on $\bar{\mathcal{Q}}_{w}$, defined as:
$\bar{Q}_{w,P}=\arg\min_{\bar{Q}\in\bar{{\cal Q}}_{w}}P\left(\bar{Q}_{P}-\bar{Q}\right)^{2}.$
Lemma 3 provides some insights and justifications on using this projection.
Lemma 3.
Let $\bar{\mathcal{Q}}_{w}=\{\theta+A\tau_{A,\beta}:\theta,\beta\}$ be the working model for the outcome regression $\bar{Q}_{P}(W,A)=E_{P}(Y\mid W,A)$, where $\theta$ is some unspecified function of $W$ and $\tau_{A,\beta}(W)=\sum_{j}\beta(j)\phi_{j}(W)$, a linear combination of basis functions. Let $\mathcal{T}_{A,w}=\{\sum_{j}\beta(j)\phi_{j}(W):\beta\}$ be the working model for the CATE function $\tau_{A,P}(W)=E_{P}(Y\mid W,A=1)-E_{P}(Y\mid W,A=0)$. Let $\bar{Q}_{w,P}=\arg\min_{\bar{Q}\in\bar{{\cal Q}}_{w}}P(\bar{Q}_{P}-\bar{Q})^{2}$ be the squared-error projection of $\bar{Q}_{P}$ onto the working model $\bar{\mathcal{Q}}_{w}$. We have that $\bar{Q}_{w,P}=E_{P}(Y\mid W)+(A-g_{P}(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}(W)$, where
$\beta_{P}=\arg\min_{\beta}E_{P}g_{P}(1-g_{P})(1\mid W)\left(\tau_{A,P}(W)-\sum_{j}\beta(j)\phi_{j}(W)\right)^{2}.$
Thus, $\sum_{j}\beta_{P}(j)\phi_{j}$ represents the projection of $\tau_{A,P}$ onto the working model $\mathcal{T}_{A,w}=\{\sum_{j}\beta(j)\phi_{j}:\beta\}$ with respect to a weighted $L^{2}$-norm using weights $g_{P}(1-g_{P})(1\mid W)$. We also have
$\beta_{P}=\arg\min_{\beta}E_{P}\left(\bar{Q}_{P}-(A-g_{P}(1\mid W))\sum_{j}\beta(j)\phi_{j}(W)\right)^{2},$
and thereby
$\beta_{P}=\arg\min_{\beta}E_{P}\left(Y-(A-g_{P}(1\mid W))\sum_{j}\beta(j)\phi_{j}(W)\right)^{2}.$
Here, one could replace $Y$ by $Y-E_{P}(Y\mid W)$. This also shows that
$\tau_{A,P}=\arg\min_{\tau\in\mathcal{T}_{A,w}}E_{P}[Y-E_{P}(Y\mid W)-(A-g_{P}(1\mid W))\tau]^{2}.$
Therefore, we can use the following loss function for learning $\tau_{A,P}$:
$L_{\theta_{P},g_{P}}(\tau)=[Y-\theta_{P}(W)-(A-g_{P}(1\mid W))\tau]^{2},$
where this loss function is indexed by the nuisance parameters $\theta_{P}(W)=E_{P}(Y\mid W)$ and $g_{P}(1\mid W)=P(A=1\mid W)$. This loss function is double robust with respect to misspecification of $\theta_{P}$ and $g_{P}$ in the sense that $\tau_{A,P}=\arg\min_{\tau\in\mathcal{T}_{A,w}}E_{P}L_{\theta,g}(\tau)$ if either $g=g_{P}$ or $\theta=\theta_{P}$.
The proof is provided in Appendix A. Interestingly, according to Lemma 3, the squared-error projection of the outcome regression $\bar{Q}_{P}$ onto the semi-parametric regression working model $\bar{\mathcal{Q}}_{w}$ corresponds to a weighted squared-error projection of the CATE function $\tau_{A,P}$ onto the corresponding working model $\mathcal{T}_{A,w}$, where the weights are $g_{P}(1-g_{P})(1\mid W)$ [adml]. As demonstrated in the efficient influence function derived in the following subsection, these weights contribute to stabilizing the variance estimate, explaining why we recommended this particular projection.
Since the true CATE function is unknown, Lemma 3 also provides a valid and practical loss function for learning CATE. This loss function is commonly referred to as the $R$-loss [nie_quasi_2021]. We could fit a relaxed highly adaptive lasso (relaxed-HAL) paired with this loss function with nuisance estimators $\theta_{n}$ for $\theta_{0}$ and $g_{n}$ for $g_{0}$ [vdl_efficient_2023]. Specifically, given a rich linear model $\{\tau_{A,\beta}=\sum_{j}\beta(j)\phi_{j}:\beta\}$ with a large set of 0th-order HAL spline basis functions $\{\phi_{j}:j\}$, we compute the lasso-estimator
$\beta_{n}=\arg\min_{\beta,\parallel\beta\parallel_{1}\leq C_{n}}\sum_{i=1}^{n}\left(Y_{i}-\theta_{n}(W_{i})-(A_{i}-g_{n}(1\mid W_{i}))\sum_{j}\beta(j)\phi_{j}(W_{i})\right)^{2},$
where $C_{n}$ is an upper bound on the sectional variation norm that equals the $L_{1}$-norm of the $\beta$-coefficients [benkeser_hal_2016]. One may also use higher-order spline HAL-MLEs as analyzed and presented in detail in [vdl_higher_2023]. The working model for $\tau_{S,P}$ is then given by ${\cal T}_{A,w}=\{\sum_{j,\beta_{n,j}\not=0}\beta(j)\phi_{j}:\beta\}$, i.e. all linear combinations of the basis functions with non-zero coefficients. Finally, we use ordinary least squares (OLS) to obtain a maximum likelihood estimator (MLE) $\beta_{n}^{\star}$ for $\beta$ within the learned working model.
4.2 Canonical gradient of $\tilde{\Psi}_{\mathcal{M}_{A,w}}$ at $P$
The working model $\mathcal{T}_{A,w}$ implies a semiparametric working model $\mathcal{M}_{A,w}$ for $P_{0}$. Let $\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{0})\equiv\tilde{\Psi}(\Pi_{\mathcal{M}_{A,w}}(P_{0}))$ be the corresponding projection estimand, where $\Pi_{\mathcal{M}_{A,w}}(P_{0})$ is the projection of $P_{0}$ onto the working model $\mathcal{M}_{A,w}$. The next step of A-TMLE involves deriving the canonical gradient (efficient influence function) of the projection parameter. First, we find the canonical gradient of the $\beta_{P}$-component.
Lemma 4.
Let $\beta_{P}$ be defined as in Lemma 3 on a nonparametric model. The canonical gradient of $\beta$ at $P$ is given by
$D_{\beta,P}^{r}=I_{P}^{-1}(A-g_{P}(1\mid W)){\bf\phi}(W)\left(Y-\theta_{P}(W)-(A-g_{P}(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}(W)\right),$
where $I_{P}=E_{P}g_{P}(1-g_{P})(1\mid W){\bf\phi}{\bf\phi}^{\top}(W)$. By Lemma 3, this can also be written as
$D_{\beta,P}^{r}=I_{P}^{-1}(A-g_{P}(1\mid W)){\bf\phi}(W)(Y-\bar{Q}_{w,P}(W,A)).$
The proof can be found in Appendix A. The next lemma presents the canonical gradient of the projection parameter $\tilde{\Psi}_{\mathcal{M}_{A,w}}$ at $P$.
Lemma 5.
Let the projection parameter $\tilde{\Psi}_{{\cal M}_{A,w}}(P)=E_{P}[\tau_{A,\beta_{P}}(W)]=E_{P}(\bar{Q}_{w,P}(W,1)-\bar{Q}_{w,P}(W,0))$, while the nonparametrically defined parameter $\tilde{\Psi}(P)=E_{P}[\tau_{A,P}(W)]=E_{P}(\bar{Q}_{P}(W,1)-\bar{Q}_{P}(W,0))$, where $\beta_{P}=\arg\min_{\beta}E_{P}g_{P}(1-g_{P})(1\mid W)(\tau_{A,P}(W)-\tau_{A,\beta}(W))^{2}$, and $\tau_{A,\beta}(W)=\sum_{j}\beta(j)\phi_{j}(W)$. We have that the canonical gradient of $\tilde{\Psi}_{{\cal M}_{A,w}}$ at $P$ is given by:
$D_{{\cal M}_{A,w},\tilde{\Psi},P}=\tau_{A,\beta_{P}}(W)-E_{P}[\tau_{A,\beta_{P}}(W)]+\sum_{j}D^{r}_{\beta,P,j}E_{P}\phi_{j}(W),$
where
$D_{\beta,P}^{r}=I_{P}^{-1}(A-g_{P}(1\mid W)){\bf\phi}(W)(Y-\bar{Q}_{w,P}(W,A)),$
with $D_{\beta,P,j}^{r}$ denoting the $j$-th component of $D_{\beta,P}^{r}$.
4.3 TMLE of the projection parameter $\tilde{\Psi}_{{\cal M}_{A,w}}$
To construct a TMLE for the pooled-ATE projection parameter, we first obtain nuisance estimators $g_{n}$ and $\theta_{n}$ of $g_{0}$ and $\theta_{0}$, respectively. If we use a relaxed-HAL for learning the working model for the conditional effect of $A$ on $Y$, then we have $\beta_{n}^{\star}=\arg\min_{\beta}P_{n}L_{\theta_{n},g_{n}}(\tau_{A,\beta})$, which solves the score equation for $\beta$, $P_{n}D^{r}_{\beta,\theta_{n},g_{n},\beta_{n}^{\star}}=0$. We estimate the marginal distribution of $W$, i.e., $P_{W,0}$, with the empirical measure of $W$ given by $P_{W,n}$. We have then solved $P_{n}D_{{\cal M}_{A,w},\tilde{\Psi},\beta_{n}^{\star},\theta_{n},g_{n},P_{W,n}}=0$. The TMLE is now the plug-in estimator $\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{n}^{\star})\equiv\tilde{\Psi}_{{\cal M}_{A,w}}(P_{W,n},\beta_{n}^{\star})$. We refer interested readers to Section 7 for a more detailed description on the implementation.